List all the term in ascending order of the finite sequence
An=7n-1 for 1 <=n <=4
There are 4 terms
Please help me with this to explain how it works! Thank you

Answer by Edwin McCravy(20056)

An = 7n-1 for 1 ≦ n ≦ 4

is short for these instructions:

1. Substitute 1 for n into 7n-1 and simplify what you get.
2. Write that down with a comma after it.
3. Substitute 2 for n into 7n-1 and simplify what you get.
4. Write that down next to the comma that you wrote in step 2, and put a comma
   after that.
5. Substitute 3 for n into 7n-1 and simplify what you get.
6. Write that down next to the comma that you wrote in step 4, and put a comma
   after that.
7. Substitute 4 for n into 7n-1 and simplify what you get.
8. Write that down next to the comma that you wrote in step 6, but don't put a
   comma after that because you are done.


Now let's do that:

1. Substitute 1 for n into 7n-1 and simplify what you get.

   7(1)-1 = 7-1 = 6

2. Write that down with a comma after it.

   6,

3. Substitute 2 for n into 7n-1 and simplify what you get.

   7(2)-1 = 14-1 = 13


4. Write that down next to the comma that you wrote in step 2, and put a comma
   after that.

   6, 13,
   
5. Substitute 3 for n into 7n-1 and simplify what you get.

   7(3)-1 = 21-1 = 20

6. Write that down next to the comma that you wrote in step 4, and put a comma
   after that.

   6, 13, 20,

7. Substitute 4 for n into 7n-1 and simplify what you get.

   7(4)-1 = 28-1 = 27

8. Write that down next to the comma that you wrote in step 6, but don't put a
   comma after that because you are done.

   6, 13, 20, 27

That list is the answer.
------------------------

Note that:
A1 = 6, A2 = 13, A3 = 20, A4 = 27.

The 1 ≦ n ≦ 4 tells you to substitute for n all the integers from 
1 to 4, inclusive, into An = 7n-1.

Edwin